hello trying to help my daughter with solving equations by elimination practice can you show me how to solve this so i can help her with these
3x-2y=-11
3x-3y=-9
and the other one is
-3x-y=-10
3x-9y=0

Answer by scott8148(6628)

you need to get terms with the same coefficient to subtract
or get terms with the same numerical coefficient, but with the opposite sign; to add
either technique will result in the elimination of the variable

subtracting the 1st two equations (to eliminate x) ___ 3x - 2y - (3x - 3y) = -11 - (-9)
___ y = -2

adding the 2nd two equations (again, to eliminate x) ___ -3x - y + (3x - 9y) = -10 + 0
___ -10y = -10

after you solve for one of the variables, you need to substitute back to find the other (eliminated) variable


you may have to multiply one or both equations by a constant to adjust the coefficients for elimination